OpenStudy (anonymous):

Urgent!!!!!! solve for x : log^{2}_{1/2} 4x + log_{2} x^2/8 = 8

OpenStudy (anonymous):

\[\log^{2}_{1/2} 4x + \log_{2} x^2/8 = 8 \]

OpenStudy (anonymous):

does log^2 even exist?

OpenStudy (anonymous):

i don't know, its in my HW

OpenStudy (anonymous):

check the statement again?

OpenStudy (anonymous):

its correct

OpenStudy (shubhamsrg):

and can you please re-post the ques,,using brackets this time..

OpenStudy (anonymous):

\[\log^{2}_{1/2} 4x + \log_{2} x^2/8 = 8 \]

OpenStudy (anonymous):

wait i ll try to solve

OpenStudy (shubhamsrg):

break log(4x) = log 4 + log x since base is 1/2 here,,we have logx -1/2= -1/log x -1/2 where base has been converted to 2 expanding other term,,we have 2 logx -3 ultimately,,on simplifying,,we will have a cubic eqn..hmmn..maybe any other mthod..?

myininaya (myininaya):

hey so @shubham.bagrecha is the way you have your problem exactly the way it appears in your book?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In my book there are 3 such questions.

OpenStudy (anonymous):

pls anyone help!!!!!!!!

myininaya (myininaya):

\[(\log_\frac{1}{2}(4x))^2+\log_2(\frac{x^2}{8})=8\] So we know the rules log(ab)=log(a)+log(b) and log(a/b)=log(a)-log(b) So using those we get \[(\log_\frac{1}{2}(4)+\log_\frac{1}{2}(x))^2+\log_2(x^2)-\log_2(8)=8\] I'm gonna use this rule log(a^r)=r log(a) \[(\log_\frac{1}{2}(4))^2+2 \cdot \log_\frac{1}{2}(4) \log_\frac{1}{2}(x)+(\log_\frac{1}{2}(x))^2+2 \log_2(x)-\log_2(8)=8\] So hmm... this looked like it didn't help

myininaya (myininaya):

What class are you in school?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

and you?

myininaya (myininaya):

I think you are looking at your problem wrong. This is like a beginning algebra class right? What book are you looking at?

OpenStudy (anonymous):

this book don't have any publisher or any author.

OpenStudy (shubhamsrg):

CBSE? NCERT ?

OpenStudy (anonymous):

no

OpenStudy (shubhamsrg):

hmm..buddy it'll form a cubic eqn,,which altough,,is solvable but i dont guess its what you'll be looking for..

myininaya (myininaya):

I'm going to take a wild guess This is a 11 grade problem Do you mean for log parts to have the same base Is this the problem: \[\log_2(4x)^\frac{1}{4}+\log_2(\frac{x^2}{8})=8\] ?

myininaya (myininaya):

\[\log_2(4x)^\frac{1}{2}+\log_2(\frac{x^2}{8})=8 ?\]

OpenStudy (anonymous):

no i've asked the right question.

OpenStudy (anonymous):

it has a log-square

OpenStudy (anonymous):

ok thanks all for whatever support you've given.

myininaya (myininaya):

Ok what I think I might have something

myininaya (myininaya):

\[\log_\frac{1}{2}(u)=\frac{\ln(u)}{\ln(\frac{1}{2})} \text{ by change of base}\] \[=\frac{\ln(u)}{\ln(1)-\ln(2)}=\frac{\ln(u)}{0-\ln(2)}=\frac{\ln(u)}{-\ln(2)}=-\frac{\ln(u)}{\ln(2)}\] \[=-\log_2(u) \text{ by change of base again }\] So we have \[(-\log_2(4x))^2+\log_2(\frac{x^2}{8})=8\] But we still have that log that is squared

myininaya (myininaya):

\[(\log_2(4x))^2+\log_2(\frac{x^2}{8})=8\]

myininaya (myininaya):

but now i'm stuck here still thinking

OpenStudy (shubhamsrg):

ohh damn..i used the same thing @myininaya ,,but i did wrong in 1 step.. you're right,,it'll be a quad eqn now now expanding..hmm

myininaya (myininaya):

Oh!

myininaya (myininaya):

Going to use log(ab)=log(a)+log(b) \[(\log_2(4)+\log_2(x))^2+\log_2(\frac{x^2}{8})=8\] using log(a/b)=log(a)-log(b) \[(\log_2(4))^2+2 \cdot \log_2(4) \log_2(x)+(\log_2(x))^2+\log_2(x^2)-\log_2(8)=8\] I did the above using (a+b)^2=a^2+2ab+b^2 now using log(a^r)=r log(a) We have \[(\log_2(4))^2+2 \cdot \log_2(4) \log_2(x)+(\log_2(x))^2+2 \log_2(x)-\log_2(8)=8\] Now gathering like terms and subtracting 8 on both sides we have \[(\log_2(x))^2+2\log_2(x)+2 \log_2(4) \log_2(x)-\log_2(8)+(\log_2(4))^2-8=0\]

myininaya (myininaya):

\[\log_2(4)=2\] so we have...

myininaya (myininaya):

and \[\log_2(8)=3 \] so we have... \[(\log_2(x))^2+2\log_2(x)+2 (2) \log_2(x)-3+(2)^2-8=0\]

myininaya (myininaya):

So combine those constants and multiply a little

OpenStudy (shubhamsrg):

alright @myininaya you did an amazing job,,now let him try a bit also :)

myininaya (myininaya):

\[(\log_2(x))^2+2\log_2(x)+4 \log_2(x)-7=0 \] Of all his now ... :)

myininaya (myininaya):

ok*